Given that the cosine of an angle is cos (∅) = 1/2°, what height of the triangle formed by this? ( The terminal point for this angle is (1,y) solve for y ).


1
Expert's answer
2022-02-24T05:50:39-0500

"y=\sin (\theta)=\sqrt{1-\cos^2(\theta)}=\sqrt{1-\frac{1}{4}}=\frac{\sqrt{3}}{2}."